OpenStudy (anonymous):

Solve the equation for x: (4x - 6)(2x - 5) = 0 a {1.5, 2.5} b {0, 2.5} c {5, 6} d no solution

OpenStudy (anonymous):

Hey.

OpenStudy (anonymous):

Do you know the FOIL method?

OpenStudy (anonymous):

Yea

hero (hero):

You don't need to FOIL to solve this. But rather, you use zero product property.

OpenStudy (anonymous):

That's all you have to do. Or that... he's probably smarter.

hero (hero):

All you do is solve 4x - 6 = 0 2x - 5 = 0

OpenStudy (anonymous):

i dont know how to solve that

hero (hero):

Can you solve 4x - 6 = 0

hero (hero):

@gabbie96

OpenStudy (anonymous):

Hero lets make a deal... I'll help her finish this if you go solve my problem. <3 http://openstudy.com/study#/updates/5239119de4b0af32a07a8f43

OpenStudy (anonymous):

lol no what do i gotta do to solve it ?

hero (hero):

Add 6 to both sides, then divide both sides by 4

OpenStudy (anonymous):

ohhh ok ok

hero (hero):

@gabbie96, what did you get for it?

OpenStudy (anonymous):

@Hero -6 ?

hero (hero):

What did you do with the 4?

OpenStudy (anonymous):

wait, im confused lol

OpenStudy (anonymous):

im only doing (4x-6) right ?

OpenStudy (anonymous):

@Hero

hero (hero):

We were doing one at a time, but we have to solve both eventually.

OpenStudy (anonymous):

i got 2.5 @Hero

hero (hero):

For which one?

OpenStudy (anonymous):

(4x-6) @Hero

hero (hero):

I don't believe you because that's not what I got for it.

hero (hero):

Add 6 to both sides, then divide both sides by 4

hero (hero):

@gabbie96

hero (hero):

Can you solve 4x - 6 = 0

hero (hero):

All you do is solve 4x - 6 = 0 2x - 5 = 0

hero (hero):

You don't need to FOIL to solve this. But rather, you use zero product property.

OpenStudy (anonymous):

what did you get ? @Hero

hero (hero):

I'll tell you in a minute, but first show me how you got 2.5

OpenStudy (anonymous):

4x - 6 +6 +6 = 10/4 = 2.5

OpenStudy (anonymous):

i think i did it wrong @Hero

hero (hero):

4x - 6 = 0 Add 6 to both sides: 4x = 6 Divide both sides by 4: x = 6/4 Simplify: x = 3/2 x = 1.5

hero (hero):

Try to solve the other one.

OpenStudy (anonymous):

ok i got 2.5

hero (hero):

Where are your steps? You didn't show them

hero (hero):

How do I know you did it right?

OpenStudy (anonymous):

oh i wrote them on my paper lol. 2x - 5 = 0 2x = 5 5/2 = 2.5

OpenStudy (anonymous):

did i do it right ?

hero (hero):

You did everything right except you forgot to express what x equals. Your steps should look like this: 2x - 5 = 0 2x = 5 x = 5/2 x = 2.5

OpenStudy (anonymous):

ohh ok ok. thanks